OpenStudy (anonymous):

Prove: sin^4x – cos^4x = sin^2x – cos^2x

Parth (parthkohli):

\[\sin(4x) - \cos(4x) = \sin^2(x) - \cos^2(x)\]?

OpenStudy (anonymous):

right

Parth (parthkohli):

Shouldn't it be \(\sin^4(x) - \cos^4(x) = \sin^2(x) - \cos^2(x)\)?

Parth (parthkohli):

Oh. OK

Parth (parthkohli):

\[\sin^4(x) - \cos^4(x) = (\sin^2(x))^2 - (\cos^2(x))^2\]Right?

Parth (parthkohli):

Hello...

OpenStudy (whpalmer4):

@ParthKohli yes, that's an identity...

Parth (parthkohli):

@isuckatmath9999 Ping

OpenStudy (anonymous):

sorry i passed out =/

Parth (parthkohli):

Hi again

Parth (parthkohli):

So \(\sin^4 (x) - \cos^4(x) = (\sin^2(x))^2 - (\cos^2(x))^2\), correct?

OpenStudy (anonymous):

why did you add on random exponents?

Parth (parthkohli):

Because I can.

Parth (parthkohli):

And it'd make my work easier...

OpenStudy (anonymous):

but then it wouldn't help me

Parth (parthkohli):

It would. Have patience :-)

OpenStudy (anonymous):

I'm on a time schedule. I don't have time to be patient I'm like 8 assignments behind >.>

OpenStudy (anonymous):

and apparently I can only do like 3 per day. So i'm fudged.

Parth (parthkohli):

But do you agree with that statement there?